# A note on generation and descent for derived categories of noncommutative schemes

### v3

@@ 3.2. Results
The hypothesus on $X$ tells us the Krull dimension of any nonempty open affine subscheme of $X$ coincides with that of $X$, cf. [Sta23, Tag 0DRT]. The regular locus of $X$ is nonempty and open, so we can find an open affine immersion $i\colon U\to X$ where $U$ is nonempty. The locus of points $p$ where $\mathcal{A}_{p}$ finitely builds $\mathcal{O}_{X,p}$ in at most one cone in $D^{b}_{\operatorname{coh}}(\mathcal{O}_{X,p})$ is open in $U$, cf. [Let21, Proposition 3.5]. Choose a basic affine open $V=\operatorname{Spec}(R)$ in this locus of points, and note that $\mathcal{A}_{p}$ finitely builds $\mathcal{O}_{X,p}$ in at most one cone in $D^{b}_{\operatorname{coh}}(\mathcal{O}_{X,p})$ for each $p\in V$. Let $j\colon V\to X$ denote the open affine immersion. Then $j^{\ast}\mathcal{A}$ finitely builds $\mathcal{O}_{V}$ in at most one cone in $D^{b}_{\operatorname{coh}}(V)$, cf. [Let21, Corollary 3.4]. Therefore, the desired claim follows immediately by Proposition 3.2.13. ∎
4. Examples
This section presents a compilation of intriguing examples exhibiting instances where Theorem 3.1 can be observed in both algebraic and geometric contexts. It is important to note that this compilation is not exhaustive and represents only a partial exploration of possible applications.
4.1. Algebras of finite type
Let $k$ be a field, and suppose $S$ is a Noetherian ring whose center $R:=Z(S)$ is of finite type over $k$. Assume that $S$ is finite over $R$. Example 3.1.7 tells us that $D^{b}_{\operatorname{coh}}(S)$ admits a classical generator. Note that $D^{b}_{\operatorname{coh}}(S)$ is considered smooth over $k$ in the sense of dg-categories, as discussed in Theorem 5.1 and Definition 2.3 of [ELS20]. From Lemma 3.5 and Lemma 3.6 in [Lun10], it follows that $D^{b}_{\operatorname{coh}}(S)$ admits a strong generator. Further details can be found in [ELS20, Remark 2.6].
4.2. Noncommutative resolutions
The forthcoming examples pertain to noncommutative commutative resolutions. This concept has a rich history, and although not exhaustively listed, key references include [VdB04, SVdB08, ŠVdB17, FMS19, DITV15, IW13]. For our discussion below, we adopt a moderately broad notion introduced in [DFI15]. Consider a commutative Noetherian ring $R$. A noncommutative resolution (NCR), as defined in [DFI15, Definition 2.5], is a finitely generated $R$-module $M$ such that $\operatorname{Supp}(M)=\operatorname{Spec}(R)$ and $\operatorname{End}_{R}(M)$ has finite global dimension. Moreover, if $R\in\operatorname{add}(M)$, then $M$ is termed a generator NCR. It is important to compare the differences in definition for an NCR in the literature. For instance, [DITV15, Definition 1.1] requires the $R$-module to be faithful, whereas [DFI15, Definition 2.5] only requires $M$ to have full support. Moreover, it is of interest to compare these examples to statement such as [IT16, Corollary 6.8] regarding properties on quotients by minimal prime ideals.
Remark 4.2.1.
Suppose $R$ is a commutative Noetherian ring. If $R$ has a generator NCR given by a finitely generated $R$-module $M$, then the Rouquier dimension of $D^{b}_{\operatorname{coh}}(R)$ is at most the global dimension of $\operatorname{End}_{R}(M)$. This follows from Corollary 3.2.
Example 4.2.2.
Any complete local reduced Noetherian ring $R$ of Krull dimension one admits a generator an NCR [DFI15, Example 2.7]. Notably, the existence of a generator NCR implies that the endomorphism ring of $R$ must have global dimension one, which, in turn, implies $R$ is a regular ring [DFI15, Proposition 3.2]. Furthermore, when $R$ is complete, it is possible to construct the generator NCR in such a way that the global dimension of its corresponding endomorphism ring is at most the multiplicity of $R$ [Leu07, Proposition 1.5].
The following statement illustrates a construction for a module $M\in\operatorname{mod}\Lambda$ satisfying the property that $\Lambda$ belongs to $\operatorname{add}_{\Lambda}(M)$ and the endomorphism ring $\operatorname{End}_{\Lambda}(M)$ has a finite global dimension. Recall a Noether algebra is a Noetherian ring $\Lambda$ that is finite over its center $\Lambda^{c}$. We denote a syzygy of module $M$ in $\operatorname{mod}\Lambda$ by $\Omega_{\Lambda}^{1}(M)$, and note that higher syzygies can be defined inductively.
Lemma 4.2.3.
Consider a Noether algebra $\Lambda$. Let $\mathcal{X}$ be a subcategory of $\operatorname{mod}\Lambda$ that is closed under finite direct sums. For any integer $n\geq 0$ and $r\geq 0$, if $M$ is an object in $\mathcal{X}$, then the following assertions hold:
  1. (1)
    Assume that $\mathcal{X}$ is closed under $r$-kernels as defined in [MT22, Definition 3.1]. For any $X\in\mathcal{X}$, if there exists an exact sequence
    $0\to M_{n}\to\cdots\to M_{0}\to X\to 0$
    with $M_{i}\in\operatorname{add}_{\Lambda}(M)$ such that the induced sequence
    $0\to\operatorname{Hom}_{\Lambda}(M,M_{n})\to\cdots\to\operatorname{Hom}_{\Lambda}(M,M_{0})\to\operatorname{Hom}_{\Lambda}(M,X)\to 0$
    is also exact, then the global dimension of $\operatorname{End}_{\Lambda}(M)$ is at most $n+r+3$.
  2. (2)
    Assume that $\mathcal{X}$ is closed under extensions. For any $N\in\Omega_{\Lambda}^{1}(\mathcal{X})$ and $L\in\mathcal{X}$, if there exists a short exact sequence
    $0\to M\to N\to L\to 0,$
    then $M\in\Omega_{\Lambda}^{1}(\mathcal{X})$. Moreover, if $\Omega_{\Lambda}^{1}(\mathcal{X})\subseteq\mathcal{X}$, then $\Omega_{\Lambda}^{1}(\mathcal{X})$ is closed under direct summands.
  3. (3)
    Assume that $\mathcal{X}$ is closed under extensions, $r$-kernels, and $\Lambda\in\mathcal{X}$. If $\Omega_{\Lambda}^{1}(\mathcal{X})\subseteq\operatorname{add}_{\Lambda}(M)$ for some $M\in\Omega_{\Lambda}^{1}(\mathcal{X})$, then $M$ is a generator NCR, and the global dimension of $\operatorname{End}_{\Lambda}(M)$ is at most $r+4$.
Proof.
[…]
$0\to X\to N_{r-1}\to\cdots\to N_{0}\to L_{1}\to L_{0},$
where $X\in\operatorname{mod}\Lambda$. By [MT22, Remark 3.2(4)], $\operatorname{add}_{\Lambda}(M)\subseteq\mathcal{X}$, implying $N_{i}\in\mathcal{X}$ for each $1\leq i\leq r$ and $L_{1},L_{0}\in\mathcal{X}$. As $\mathcal{X}$ is closed under $r$-kernels, we deduce that $X\in\mathcal{X}$. Given our hypothesis, there exists an exact sequence
$0\to M_{n}\to\cdots\to M_{0}\to X\to 0$
with $M_{i}\in\operatorname{add}_{\Lambda}(M)$, ensuring that
$0\to\operatorname{Hom}_{\Lambda}(M,M_{n})\to\cdots\to\operatorname{Hom}_{\Lambda}(M,M_{0})\to\operatorname{Hom}_{\Lambda}(M,X)\to 0$
is also exact. Splicing with the exact sequence
$0\to\operatorname{Hom}_{\Lambda}(M,X)\to\operatorname{Hom}_{\Lambda}(M,N_{r-1})\to\cdots\to\operatorname{Hom}_{\Lambda}(M,N_{0})\to\operatorname{Hom}_{\Lambda}(M,N)\to 0,$
we obtain the following
$0\to$ $\operatorname{Hom}_{\Lambda}(M,M_{n})\to\cdots\to\operatorname{Hom}_{\Lambda}(M,M_{0})$
$\to\operatorname{Hom}_{\Lambda}(M,N_{r-1})\to\cdots\to\operatorname{Hom}_{\Lambda}(M,N_{0})\to\operatorname{Hom}_{\Lambda}(M,N)\to 0$
where each $M_{i},N_{j}\in\operatorname{add}_{\Lambda}(M)$, hence each $\operatorname{Hom}_{\Lambda}(M,M_{i}),\operatorname{Hom}_{\Lambda}(M,_{j})\in\operatorname{proj}T$. Since applying $\operatorname{Hom}_{\Lambda}(M,-)$ to the exact sequence $0\to N\to L_{1}\to L_{0}$ also gives us an exact sequence
$0\to\operatorname{Hom}_{\Lambda}(M,N)\to P_{1}\to P_{0}\to Y\to 0,$
we conclude $\operatorname{pdim}_{T}Y\leq n+r+3$. Since $Y\in\operatorname{mod}T$ is arbitrary, this finishes the claim.
(2) The validity of the latter part of the assertion is evident from the former, achieved by expressing the split short exact sequence corresponding into a direct sum. To establish the first part of the claim, under the given hypothesis, consider the exact sequence $0\to N\to P\to N^{\prime}\to 0$ for some $P\in\operatorname{proj}\Lambda$ and $N^{\prime}\in\mathcal{X}$. The desired conclusion can then be derived from the following pushout diagram:
[…]
remembering that $U\in\mathcal{X}$ because the subcategory $\mathcal{X}$ is extension-closed.
(3) We establish $\Omega_{\Lambda}^{1}(\mathcal{X})\subseteq(\mathcal{X})$ as per [MT22, Remark 3.2(4)]. Consequently, by claim (2), $\Omega_{\Lambda}^{1}(\mathcal{X})$ is closed under direct summands, leading to $\Omega_{\Lambda}^{1}(\mathcal{X})=\operatorname{add}_{\Lambda}(M)$. In particular, this implies that $M$ is a generator NCR. Let $T=\operatorname{End}_{\Lambda}(M)$, and choose $X\in\mathcal{X}$. We know that $\operatorname{Hom}_{\Lambda}(M,X)\in\operatorname{mod}T$. Consider a map $f\colon M_{0}\to X$, where $M_{0}\in\operatorname{add}_{\Lambda}(M)$, and observe that the sequence
$\operatorname{Hom}_{\Lambda}(M,M_{0})\xrightarrow{\operatorname{Hom}_{\Lambda}(M,f)}\operatorname{Hom}_{\Lambda}(M,X)\to 0$
forms a $T$-projective cover of $\operatorname{Hom}_{\Lambda}(M,X)$. Since $M$ is a generator NCR, it follows that $f$ is surjective. As $M_{0}\in\Omega_{\Lambda}^{1}(\mathcal{X})$, we deduce that $\ker f\in\Omega_{\Lambda}^{1}(\mathcal{X})$ by claim (2). Consequently, the sequence
$\sigma\colon 0\to\ker f\to M_{0}\to X\to 0$
is exact, where $\ker f,M_{0}\in\operatorname{add}_{\Lambda}(M)$, and $\operatorname{Hom}_{\Lambda}(M,\sigma)$ is also exact. As $X\in\mathcal{X}$ was chosen arbitrarily, we have demonstrated that the hypotheses of property (1) are satisfied with $n=1$. Therefore, the global dimension of $\operatorname{End}_{\Lambda}(M)$ is at most $r+4$. ∎
Remark 4.2.4.
To apply the above construction in a down-to-earth scenario, consider a Cohen-Macaulay ring $\Lambda$ with Krull dimension $r+1$. Note $\Lambda$ is commutative. In this case, let $\mathcal{X}:=\operatorname{CM}(\Lambda)$, which includes $\Lambda$ and is closed under extensions and $r$-kernels. Notably, if $R$ is a normal domain with Krull dimension two and $\operatorname{Spec}(R)$ has rational singularities, then, as per [IW10, Theorem 3.6(a)] and the above construction, $R$ possesses a generator NCR. Conversely, for a local, excellent, Henselian domain $R$ with an algebraically closed residue field and admitting a generator NCR, the conclusion holds that $\operatorname{Spec}(R)$ has rational singularities, cf. [DITV15, Corollary 3.3].
Example 4.2.5.
Let $R$ be a local Cohen-Macaulay ring of finite Cohen-Macaulay representation type. There exists a generator NCR $M\in\operatorname{mod}R$ such that the global dimension of $\operatorname{End}_{R}(M)$ is at most $\max\{2,\dim R\}$. This is [Leu07, Theorem 2.1].
4.3. Categorical resolutions
There is a homological construction, called categorical resolutions, that emulates the functorial behavior observed in the derived category of a variety with rational singularities. A substantial body of work has been devoted to this pursuit, and the following list of key references is not complete: [KL15, Lun10, Kuz07, BBF16, DD23, Lun12, KS22, Abu16]. Let $X$ be a separated Noetherian scheme of finite type over a field. A categorical resolution of $X$ is a smooth cocomplete compactly generated triangulated category $\mathcal{T}$ with a pair of adjoint functors $\pi_{\ast}\colon\mathcal{T}\to D_{\operatorname{Qcoh}}(X)$ and $\pi^{\ast}\colon D_{\operatorname{Qcoh}}(X)\to\mathcal{T}$ such that
  1. (1)
    $\pi_{\ast}\pi^{\ast}\cong 1_{D_{\operatorname{Qcoh}}(X)}$
  2. (2)
    $\pi^{\ast}$ and $\pi_{\ast}$ commute with small coproducts
  3. (3)
    $\pi^{\ast}(\mathcal{T}^{c})$ is contained in $D^{b}_{\operatorname{coh}}(X)$ where $\mathcal{T}^{c}$ is the subcategory of compact objects in $\mathcal{T}$.
There is a variant for a categorical resolution of $D^{b}_{\operatorname{coh}}(X)$, cf. [Kuz07, Definition 3.2], but this is not treated here.
Example 4.3.1.
Let $X$ be a reduced separated Noetherian one-dimensional scheme that is of finite type and projective over a field $k$. Suppose $\nu\colon\widetilde{X}\to X$ is the normalization of $X$. There exists an inductively constructed collection of $\mathcal{O}_{X}$-modules:
$\mathcal{O}_{1}:=\mathcal{O}_{X}\subseteq\mathcal{O}_{2}\subseteq\mathcal{O}_{3}\subseteq\cdots\subseteq\mathcal{O}_{n}:=\nu_{\ast}\mathcal{O}_{\widetilde{X}}$
[…]
Example 4.3.2.
Consider a separated Noetherian scheme of finite type over a field $k$. Suppose there exists an non-negative integer $n$ and ideal sheaf $\mathcal{I}$ on $X$ such that $\mathcal{I}^{n}=(0)$. Consider the coherent sheaf of $\mathcal{O}_{X}$-algebras
[…]
Assume the closed subscheme $Z$ of $S$ corresponding to the ideal $\mathcal{I}$ is smooth. There exists a semiorthogonal decomposition
$D^{b}_{\operatorname{coh}}(\mathcal{A})=\langle D^{b}_{\operatorname{coh}}(Z),\cdots,D^{b}_{\operatorname{coh}}(Z)\rangle$
where the number of components is equal to $n$. This is [KL15, Corollary 5.15]. Moreover, $D^{b}_{\operatorname{coh}}(\mathcal{A})$ is a categorical resolution of $D^{b}_{\operatorname{coh}}(Z)$, cf. [KL15, Theorem 5.23].
4.4. Homological projective duality
There is a triangulated framework, due to [Kuz06], that enables a rich comparison of the bounded derived categories of two projectively dual varieties which is called homological projective duality (HPD). Informally, HPD states the derived categories of linear sections of a smooth projective mapping to a projective space $X\to\mathbb{P}(V)$ are governed by a (noncommutative) variety $Y\to\mathbb{P}(V^{\vee})$ over the dual projective space, where $Y$ is called the HPD dual of $X$. Recently, HPD has been generalized to broader (noncommutative) contexts [Per19, KP21]. See [Kuz14, Tho18] for survey and additional background. Let $k$ be a field.
Example 4.4.1.
Let $U,V$ be vector spaces over $k$ respectively with dimensions $m,n$. Assume $n\geq m$. Set $W:=U\otimes_{k}V$. Choose an integer $11\leq r\leq m-$ 1. Define $\mathcal{L}^{r}:=\mathcal{L}^{r}_{m,n}$ to be the variety of matrices $M\colon V\to U^{\vee}$ in $\mathbb{P}(W)$ cut by the minors of size $r+1$ of the matrix of indeterminates $(x_{ij,})_{i,j=1}^{m,n}$. Consider the Grassmannian variety $\mathbb{G}:=\mathbb{G}(U,r)$ of $r$-dimensional quotient spaces of $U$. Denote the quotient bundle and tautological subbundle over $\mathbb{G}$ of respective rank $r$ and $m-r$ by $\mathcal{D}$ and $\mathcal{U}$. The Euler sequences gives us
$0\to\mathcal{U}\to U\otimes_{k}\mathcal{O}_{\mathbb{G}}\to\mathcal{D}\to 0$
[…]
Example 4.4.2.
Suppose $V$ is a $k$-vector space of odd dimension $v$. For each even number $0\leq 2q<v$, we have a Pfaffian variety $\operatorname{Pf}_{q}$ contained in $\mathbb{P}(\wedge^{2}V^{\vee})$ consisting of all $2$-forms on $V$ whose rank is at most $2q$. This variety is not a complete intersection and usually is singular. The projective dual of $\operatorname{Pf}_{q}$ is another Pfaffian variety whose locus is the subspace of $\mathbb{P}(\wedge^{2}V)$ consisting of bivectors of rank at most $2s$ where $2s=v-1-2q$. There are noncommutative crepant resolutions for $\operatorname{Pf}_{s}$ and $\operatorname{Pf}_{q}$ respectively denoted by $A$ and $B$, see [ŠVdB17, Section 6]. Note this gives noncommutative schemes $(\operatorname{Pf}_{s},A)$ and $(\operatorname{Pf}_{q},B)$. These are homological projective duals of one another, cf. [RS19]. Let $L$ be a generic subspace of $\wedge^{2}V^{\vee}$ with dimension $sv$. Moreover, the sheaves $A,B$ restrict to give noncommutative crepant resolutions of the varieties $\operatorname{Pf}_{s}\cap\mathbb{P}(L^{\perp}),\operatorname{Pf}_{q}\cap\mathbb{P}(L)$ such that $D^{b}_{\operatorname{coh}}(A|_{\mathbb{P}(L^{\perp})})$ and $D^{b}_{\operatorname{coh}}(B|_{\mathbb{P}(L)})$ are equivalent categories. This is [RS19, Theorem 1.1]. The reader is encouraged to see [RS19] for further details; including the case of even-dimensional vector spaces.

### v4

@@ 4. Descent along structure morphism
We know that $D^{b}_{\operatorname{coh}}(\mathcal{A})$ admits a classical generator, see [ELS20, Theorem 4.15]. If $D^{b}_{\operatorname{coh}}(\mathcal{A})$ has no strong generators, then there is nothing to check, so we can assume $D^{b}_{\operatorname{coh}}(\mathcal{A})$ admits a strong generator. Suppose $G$ is a strong generator for $D^{b}_{\operatorname{coh}}(\mathcal{A})$ with minimal generation time. The regular locus of $X$ is nonempty and open, so we can find an affine open immersion $i\colon U\to X$ where $U$ is nonempty. The locus of points $p$ where both $(\pi_{\ast}G)_{p}$ and $\mathcal{O}_{X,p}$ finitely build one another in at most one cone in $D^{b}_{\operatorname{coh}}(\mathcal{O}_{X,p})$ is open in $U$, see [DL24b, Lemma 3.1] or [Let21, Proposition 3.5].
Let $V$ be an affine open contained in the intersection of $U$ and such points. [^10] This is nonempty because both objects do not vanish at the generic point. Denote by $j\colon V\to X$ the open immersion. It follows that both $j^{\ast}\pi_{\ast}G$ and $\mathcal{O}_{V}$ finitely build one another in one cone in $D^{b}_{\operatorname{coh}}(V)$, see [Let21, Corollary 3.4]. Note $V$ is contained in the regular locus of $X$. This tells us $V$ is a affine regular scheme of finite Krull dimension, so $\mathcal{O}_{V}$ is a strong generator whose generation time is equal to $\dim V$, see [Let20, Corollary 4.3.13] or [Chr98, §$\S 8.2$]. Hence, $j^{\ast}\pi_{\ast}G$ has generation time coinciding with $\dim V$ in $D^{b}_{\operatorname{coh}}(V)$.
The hypothesis on $X$ tells us the Krull dimension of any nonempty affine open subscheme of $X$ coincides with that of $X$, see [Sta23, Tag 0DRT]. This ensures that $j^{\ast}\pi_{\ast}G$ is a strong generator for $D^{b}_{\operatorname{coh}}(V)$ with generation time being $\dim X$. Choose a closed point $p$ in $V$ such that $\dim\mathcal{O}_{V,p}=\dim X$. As $V$ is regular, it follows that $\operatorname{level}^{\mathcal{O}_{V,p}}(\kappa(p))=\dim X+1$ see [Let20, Corollary 4.3.13] coupled with [Sta23, Tag 00OB]. Let $s\colon\operatorname{Spec}(\kappa(p))\to V$ be the closed immersion. There is a commutative diagram of exact functors:
${{D^{b}_{\operatorname{coh}}(\kappa(p))}}$${{D^{b}_{\operatorname{coh}}(X)}}$${{D^{b}_{\operatorname{coh}}(s^{\ast}j^{\ast}\mathcal{A})}}$${{D^{b}_{\operatorname{coh}}(j^{\ast}\mathcal{A})}.}$$\scriptstyle{s_{\ast}}$$\scriptstyle{s_{\ast}}$$\scriptstyle{\pi_{\ast}^{\prime}}$$\scriptstyle{\theta_{\ast}}$
where $\theta\colon\mathcal{O}_{\operatorname{Spec}}(\kappa(p))\to s^{\ast}j^{\ast}\mathcal{A}$ and $\pi^{\prime}\colon\mathcal{O}_{V}\to j^{\ast}\mathcal{A}$ are the structure morphisms.
We see that $\theta_{\ast}s^{\ast}j^{\ast}\mathcal{A}$ is isomorphic to a complex of the form $\bigoplus_{n\in\mathbb{Z}}\mathcal{O}_{\operatorname{Spec}}(\kappa(p))^{\oplus r_{n}}[n]$. Recall that $j^{\ast}\colon D^{b}_{\operatorname{coh}}(\mathcal{A})\to D^{b}_{\operatorname{coh}}(j^{\ast}\mathcal{A})$ is a Verdier localization, see Remark 3.5. Suppose $Q$ is an object of $D^{b}_{\operatorname{coh}}(\mathcal{A})$ such that $j^{\ast}Q$ is isomorphic to $s_{\ast}s^{\ast}j^{\ast}\mathcal{A}$ in $D^{b}_{\operatorname{coh}}(j^{\ast}\mathcal{A})$. The following diagram commutes up to a natural isomorphism, see proof of [DLM24a, Proposition 3.1]:
${{D^{b}_{\operatorname{coh}}(j^{\ast}\mathcal{A})}}$${{D^{b}_{\operatorname{coh}}(U)}}$${{D^{b}_{\operatorname{coh}}(\mathcal{A})}}$${{D^{b}_{\operatorname{coh}}(X)}.}$$\scriptstyle{\pi^{\prime}_{\ast}}$$\scriptstyle{\pi_{\ast}}$$\scriptstyle{j^{\ast}}$$\scriptstyle{j^{\ast}}$
This tells us that $\pi^{\prime}_{\ast}j^{\ast}G$ is isomorphic to $j^{\ast}\pi_{\ast}G$ in $D^{b}_{\operatorname{coh}}(V)$. Tying together these observations gives us the following string of inequalities:
$\dim D^{b}_{\operatorname{coh}}(\mathcal{A})+1$ $=\operatorname{gen.time}(G)+1$
$\geq\operatorname{level}^{G}(Q)$
$\geq\operatorname{level}^{j^{\ast}G}(j^{\ast}Q)$
$\geq\operatorname{level}^{\pi^{\prime}_{\ast}j^{\ast}G}(\pi^{\prime}_{\ast}j^{\ast}Q)$
$\geq\operatorname{level}^{j^{\ast}\pi_{\ast}G}(\pi^{\prime}_{\ast}j^{\ast}Q)$
$\geq\operatorname{level}^{\mathcal{O}_{V}}(\mathcal{O}_{\operatorname{Spec}}(\kappa(p)))$
$\geq\operatorname{level}^{\mathcal{O}_{V,p}}(\kappa(p))=\dim X+1.$
This completes the poof. ∎